OpenStudy (anonymous):

What is the slope of the line whose equation is 3x + 6y = 9? A. -3 B.-1/2 C.1/2 D.6

OpenStudy (anonymous):

solve for y \[y=(-1/2) x+3/2\]

OpenStudy (the_fizicx99):

3x + 6y = 9 -3x -3x 6y = -3x + 9 /6 /6 /6 y = -3/6x + 1.5 y = -1/2x + 1.5

OpenStudy (the_fizicx99):

@perdomo.arturo you're not suppose to give out answers. Guide.

OpenStudy (anonymous):

but how do i figure out x

OpenStudy (the_fizicx99):

What do you mean x?

OpenStudy (anonymous):

you have y= -1/2X - 1.5 how do i find x

OpenStudy (anonymous):

+1.5*

OpenStudy (the_fizicx99):

x is your slope. y = mx + b

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

@tHe_FiZiCx99 I didn't give the answer. I solved for y and assumed broklinngaylee new how to get the slope from there

OpenStudy (the_fizicx99):

What I mean is M is the slope. So -1/2x is your slope. B is they y intercept. In this case its 1.5. You would plot 1.5 on the y axis. @perdomo.arturo she's asking because she doesn't understand. So you guide her through it.

OpenStudy (anonymous):

okay but i dont have graph paper and it doesnt give me a graph on the question its through k12

OpenStudy (anonymous):

okay i get it now thank you both!!!:)

OpenStudy (the_fizicx99):

No you don't need it, they're just asking to find the slope (M) of the given equation. They put it in standard form (Ax + By = C) so you had to convert it to slope intercept form.

OpenStudy (anonymous):

okay! thank you for guiding me through this problem i understand it now!:)

OpenStudy (the_fizicx99):

Yw!

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html ,you check for way work out here